OpenStudy (anonymous):

there are 30 students in the choir at gold elementary. One- third of the students are 6th graders, and the rest are 5th graders. Of the sixth grade choir 1/5 won 1st place in the choir concert this summer. How many 6th grade students won first place in the choir concert?

OpenStudy (anonymous):

6th grade = 30*(1/3)=10 won = 10*(1/5) = 2 so 2